Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability Q..pls answer

Expert replies
by Inspired » Thu Jan 14, 2010 11:12 am
If a six sided die is rolled three times, what is the probability of getting at least one even number and at least one odd number?

1/8
1/4
1/2
3/4
7/8

pls ans with explanation
Join the discussion
Source: — Problem Solving |

by Brent@GMATPrepNow » Thu Jan 14, 2010 11:25 am
Inspired wrote:If a six sided die is rolled three times, what is the probability of getting at least one even number and at least one odd number?

1/8
1/4
1/2
3/4
7/8

pls ans with explanation
For questions involving "at least" it is often useful to use the complement.
That is: P(event A occurs) = 1 - P(event A does not occur)

In your question, what would it mean if your event (at least one even and one odd) does not occur?
It would mean we have either all odds or all evens.

So, we are looking for 1 - P(all odd or all even)

P(all odd OR all even) = P(all odd) + P(all even)

Begin with P(all odd)
What is the probability that all three numbers are odd?
Well, the probability that the first roll is odd = 3/6 = 1/2
The probability that the 2nd roll is odd = 1/2
The probability that the 3rd roll is odd = 1/2
So, P(1st is odd AND 2nd is odd AND 3rd is odd) = 1/2 x 1/2 x 1/2 = 1/8

We can perform similar calculations to see that P(all even) = 1/8

So, P(all odd OR all even) = P(all odd) + P(all even)
= 1/8 + 1/8
= 1/4

So, 1 - P(all odd or all even) = 1 -1/4 = 3/4 (D)
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Inspired » Thu Jan 14, 2010 11:35 am
thank you very much for your detailed explanation. I am now clear about AND and OR operators too.
Join the discussion